OpenStudy (anonymous):

evaluate integral of sin^(2)xcos^(4)xdx, the answer on the book is (-1/6)cos^(5)xsinx+(1/24)cos^(3)xsinx+(1/16)cosxsinx+(1/16)x+C, for some reason i got something different when i did it

OpenStudy (tkhunny):

What was your method of operation? Are you SURE your result was not equivalent?

OpenStudy (anonymous):

you're right, both answers are equivalent, i just checked, thank you

OpenStudy (tkhunny):

Good work. It is sometimes a little confusing. Remember, with that "+C", thing going on, these are IDENTICAL results: \(\sin^{2}(x) + C_{1} = \cos^{2}(x) + C_{2}\)